OpenStudy (anonymous):

What is the value of the expression? (–4)4

OpenStudy (alexandervonhumboldt2):

-4*4=-(4*4)=-16

OpenStudy (anonymous):

uhhh

OpenStudy (anonymous):

-16 because a negative times a positive is a negative Is that a better explanation? (:

OpenStudy (anonymous):

:)thanks :)

OpenStudy (anonymous):

You are quite welcome!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@PrincessHush ik i ve already done this lol

OpenStudy (phi):

do you mean \[ (-4)^4 \] or \[ (-4)4\]

OpenStudy (anonymous):

(−4)4

OpenStudy (anonymous):

first onbe

OpenStudy (anonymous):

one*

OpenStudy (phi):

with a small 4 in the upper right ?

OpenStudy (phi):

you should use an ^ to show that (-4)^4 that "little 4" is an exponent, and it is short for multiply the "big number" times itself 4 times \[ (-4)^4 = -4 \cdot -4 \cdot -4 \cdot -4 \]

OpenStudy (anonymous):

ohhh